# Strategic Tradeoffs Between Humans and AI in Multi-Agent Bargaining

### v3

@@ 5.1 Human subjects
We recruited 216 U.S.-based participants from Prolific to complete the bargaining game on Deliberate Lab (Tsai et al., 2024). [^4] IRB approval and informed consent were obtained; we applied no inclusion criteria for recruitment. Participants completed two games over a 30-minute session, which included time waiting for two others to join a live game. To incentivize strategic play, participants received the surplus value of their final chip holdings in addition to a \$4 base payment. The average total payout was \$12.24 $\pm$ \$6.12. Each participant completed two games with different chip valuations and profiles. For each of the 2-, 3-, and 4-chip game variants, 24 groups of three players played 2 games each, yielding $N=3\times 24\times 3\times 2=432$ effective participants across 144 total games. [^5] The two games are evaluated independently, as ordering and learning effects were minimal.
5.2 LLM agents
@@
We then replicated each human game with LLM agents. At each turn, an LLM agent received the same information as its human counterpart, including private chip valuations, public chip holdings, and trade history. One agent proposed a trade per round; the others evaluated it.
We developed two types of agents to evaluate the tradeoffs between simple prompting and chain-of-thought refinement. As our primary goal was to evaluate baseline performance under minimal optimization, we first constructed an out-of-box agent used simple prompts adapted from the human instructions. To assess the benefit of lightweight refinement, we also implemented a refined agent which generated candidate proposals and selected one using “type-2” reasoning, mimicking slower, more deliberate decision-making (Furniturewala et al., 2024). Each agent type was implemented across two model families for generalizability: OpenAI’s GPT-4o (OpenAI et al., 2024) and Google’s Gemini 1.5 Pro (Team et al., 2024), which were the latest full-capacity models from each family at time of writing. [^6] For robustness, we also benchmarked the latest lightweight models (Gemini 2.5 Flash, GPT-o4-mini), but they failed to execute basic surplus-maximizing trades. We therefore restrict our main analysis to full-capacity models; see Appendix F for additional context. Prompt temperature was set to 0.5, following mid-range recommendations for structured reasoning tasks that require coherence and variability (Arora et al., 2024). Prompts were scaffolded using a chain-of-thought strategy through the open-source EDSL library (Wei et al., 2022; Horton et al., 2024).
5.3 Bayesian agents
@@
We include Bayesian players as a statistical benchmark: these agents represent the traditional, hard-coded approach to negotiation, explicitly programmed to optimize within the structure of this bargaining game. Each agent i forms a joint belief about the valuations of all other players, $\mathbf{v}_{-i}$. When agent i proposes a trade $(x_{g},y_{r})$ – offering $x$ chips of type $g$ in exchange for $y$ chips of type $r$ – they choose a trade to maximize their expected payoff:
@@
$\max_{(\mathbf{x},\mathbf{y})}\sum_{\mathbf{v}_{-i}}\mathbf{1}\Bigl\{\exists\,j\neq i:\text{accept}\bigl(v_{j},x_{g},y_{r}\bigr)\Bigr\}\;\times\;\Delta u_{i}(v_{i},x_{g},y_{r})\;\times\;B_{i}\bigl(\mathbf{v}_{-i}\bigr),$ (1)
@@
where $\text{accept}\bigl(v_{j},x_{g},y_{r}\bigr)$ indicates that player j with valuation $v_{j}$ would accept the trade, $\Delta u_{i}(v_{i},x_{g},y_{r})=v_{i,r}y_{r}-v_{i,g}x_{g}$ is player $i$’s change in utility and $B_{i}(\mathbf{v}_{-i})$ is player $i$’s belief (probability) that the other players’ valuations are $\mathbf{v}_{-i}$. This agent assumes that their opponent’s acceptance decision is myopically rational, i.e. a receiver with valuation $v_{j}$ accepts if and only if they have sufficient chips for the trade and their expected utility is positive, i.e. $v_{j,g}x_{g}-v_{j,r}y_{r}>0$.
@@
All players update their priors after observing acceptance decisions. A player discards valuation states $\{v_{j}\}$ that contradict the observed event, assuming myopically rational opponents, and re-normalizes the remaining probabilities. When the trade is accepted, the proposer keeps only valuations that would be accepted in its distribution over the chosen receiver; the receiver keeps only valuations that would be accepted in its distribution over the proposer. Bystanders update priors for both accordingly. When there is no trade, all players discard valuations that conflict with the observed non-acceptance.
Figure 2: Surplus trajectories for human, LLM, and Bayesian-learning agent simulations across game complexities. Each unique game yields a blue line corresponding to the ratio of surplus achieved relative to the computed optimal allocation. Means and 95% confidence intervals are highlighted and listed in Appendix 3. All populations generated positive aggregate surplus.
6 Results
6.1 Performance
Figure 2 shows surplus trajectories by agent type and game variation. Bayesian agents reached 74% of the Pareto-efficient optimum in the two-chip game (upper right panel). Across all variations, Bayesian agents achieved the highest surplus. GPT-4o performed comparably to human participants; however, humans statistically outperformed Gemini 1.5 Pro. [^7] This may reflect the recency of such capabilities; Gemini 1.5 Pro was released three months prior to GPT-4o and represented an earlier generation of model development. Refined LLMs did not significantly impact surplus compared to out-of-the-box LLMs (Appendix E).
6.2 Procedural alignment
In addition to performance, we analyze procedural alignment, how the agents engage in the mechanics of negotiation, through two lenses: 1) observed trading patterns, and 2) regret minimization strategies.
6.2.1 Trading patterns and volume dynamics
Figure 3: Trading patterns in the 3-chip game, visualizing (i) net surplus change for the proposer and (ii) trade ratio. Accepted trades are shown in green, rejected trades in red. Marginal distributions across each axis are adjacent to the plots. The vertical dashed line marks zero net surplus (no net value created), and the horizontal solid line marks balanced exchange (1:1 ratio of chips). Figure values are provided in Table 5.
Humans and social norms.
Figure 3 plots individual trade proposals for the three-chip game for each agent type. Results for the two- and four-chip games are qualitatively similar (Appendix F.2). The vertical balanced value line indicates proposals that generated no net surplus, while the horizontal balanced ratio line corresponds to 1:1 exchanges of chips. In the left panel, we see that the human proposals cluster around the balanced ratio line: that is, humans often offer a similar number of chips as they are requesting. Most trade proposals lie to the right of the balanced value line (i.e., provide positive surplus to the proposer), though some reduce total surplus.
LLMs and concessionary exchanges.
Both LLM families propose trades near the vertical balanced net value line, yielding little total surplus with no anchoring to balanced trade ratios. A noticeable vertical tail of proposals extends above the 1:1 balanced ratio line, indicating a tendency to offer more chips than requested (even up to 5:1). Many proposals also incur a net surplus loss.
Bayesian agents and rational surplus maximization.
By construction, Bayesian agents strictly maximize surplus (evidenced by positive surplus proposals) and avoid net-loss trades (no points left of the balanced value line). Rather than the 1:1 balanced ratio exhibited in human proposals, the Bayesian offers cluster below the parity line (ratio < 1), meaning they ask for more chips than they give. Proposal rejections are high.
Trade pattern comparisons.
@@ 6.2.1 Trading patterns and volume dynamics
In sum, these trading patterns reveal clear strategic differences: humans offer trades suggesting a consideration for balanced, fair exchanges (Danz et al., 2022), LLMs adopt a “concessionary posture” with high proposal acceptance rates, and Bayesian agents adopt a value-extractive strategy (Fisher et al., 2011), asking to receive more than they give and incurring higher rejection rates. To evaluate whether any of these strategies were more optimal ex-post, we next evaluate these trades using a regret-minimization framework.
6.2.2 Strategic alignment and regret minimization
At each turn, a participant might take an action of proposing, accepting, or declining a trade. In no regret actions, no feasible later alternative would yield a better valuation for the target chip. Forced regret occurs when a higher-surplus option arises later, but the agent’s earlier decisions—such as premature trades or overcommitting inventory—prevent acting on it. Finally, unforced regret captures scenarios where a better option becomes available, yet the agent fails to take it.
To avoid speculative modeling of unobserved strategies, we limit the scope of our analysis. First, we only consider myopically rational transactions (i.e., positive surplus) for proposers, acceptors, or decliners. [^8] Bayesian agents satisfy this by design. Interpreting irrational trades, which may be driven reciprocity, signaling, or cognitive error, would require assumptions we cannot credibly test. Second, when evaluating proposer behavior, we focus on accepted trades; rejected offers may result from strategic signaling by either party or proposer misjudgment of demand. [^9] This analysis is distinct from evaluations conducted from the perspective of decliners. Third, “acceptors” include those who intended to accept a trade, not only those who were selected to trade. [^10] This occurs when both accept the trade and one is randomly chosen. This preserves the behavioral signal of willingness to accept, without conflating risk tolerance or random assignment.
Figure 4 shows an example of a player’s valuations as they trade over the course of a single game. [^11] The sum of all players’ surplus changes at each turn correspond to the blue lines in Figure 2. On turn 1, Player 3 exhibited a forced regret action as an acceptor: they accepted a trade which resulted in their having insufficient chips to accept a more profitable offer in turn 2. On turn 7, Player 2 exhibited an unforced regret action as a decliner; they declined a profitable trade and did not recover the profits later. On Turn 8, Player 2 made a no regret proposal; this was the best deal they could have made, as there were no better future trades. [^12] More detailed explanations are provided in Appendix D.

### v4

@@ 5.1. Human Subjects
We recruited 216 U.S.-based participants from Prolific to complete the bargaining game on Deliberate Lab. [^6] Institutional IRB approval and informed consent were obtained; we applied no inclusion criteria for recruitment. Each participant completed two games over a 30-minute session. To incentivize play, one game was randomly selected for payment, and participants received their surplus—the difference between their final and initial chip values—as a bonus in addition to a \$4 base payment. For example, a player whose chips increase in value from \$10 to \$14 earns a \$4 surplus bonus. The average total payout was \$12.24 $\pm$ \$6.12. Each participant completed two games with different chip valuations and profiles. For each of the 2-, 3-, and 4-chip game variants, 24 groups of three players played 2 games each with 9 turns, yielding 1,296 trade proposals across 144 games. Participants were not informed whether they were matched with the same partners across games. The two games are evaluated independently, as ordering and learning effects proved minimal. [^7] Learning effects are discussed in Appendix F.1.
5.2. LLM Agents
We replicated each human game with two types of LLM agents. First, we wanted to evaluate how LLMs perform “out of the box,” constructing a baseline agent using intentionally simple prompts adapted directly from the human instructions. Second, we wanted to see if we could improve performance by prompting the LLM to reason more deeply about the strategic context. To do this, we implemented a refined agent that first generates multiple candidate proposals and then uses a second LLM call to select among them—mimicking the slower, more deliberate “Type 2” decision-making process described by Kahneman (2011) and applied to LLMs by Furniturewala et al. (2024). We tested both prompting strategies on OpenAI’s gpt-4o-2024-05-13 (gpt-4o, hereinafter) (OpenAI et al., 2024) and Google’s gemini-1.5-pro-001 released on 2024-05-23 (Gemini 1.5 Pro) (Team et al., 2024), which were the latest full-capacity models from each family at the time of writing. [^8] We also tested (Gemini 2.5 Flash, GPT-o4-mini), but they could not reliably propose valid trades. See Appendix G for further details. Model temperature was set to 0.5.999A lower temperature risked agents repeatedly sampling a narrow subset of trades (Lambert and Calandra, 2023). A higher temperature risked invalid trades (Arora et al., 2024). Prompts were designed using chain-of-thought (Wei et al., 2022), and simulations were implemented using the open-source EDSL library (Horton et al., 2024).
@@ 5.2. LLM Agents
Importantly, the replications were identical to the human games in terms of initial conditions and gameplay. At each turn, an LLM agent received the same information as its human counterpart, including private chip valuations, public chip holdings, and trade history. One agent proposed a trade per round, and the others evaluated it.
Figure 2. Surplus trajectories across game complexities by agent type. Each blue line traces a game’s surplus as a fraction of the Pareto-optimal benchmark; black lines show average trajectories with 95% confidence bands. The dashed line marks the theoretical ceiling (ratio = 1), and each panel displays the final average at Turn 9. Final surplus values are reported in Table 3.
Surplus trajectories across game complexities by agent type. Each blue line traces a game’s surplus as a fraction of the Pareto-optimal benchmark; black lines show average trajectories with 95% confidence bands. The dashed line marks the theoretical ceiling (ratio = 1), and each panel displays the final average at Turn 9.
5.3. Bayesian Agents
@@
We include a Bayesian agent as a task-specific benchmark—an explicitly programmed agent that optimizes within the known structure of the game. We summarize the algorithm here and provide full details in Appendix A.3. Each agent $i$ forms a joint belief about the valuations of all other players, $\mathbf{v}_{-i}$. When agent $i$ proposes a trade $(x_{g},y_{r})$ – offering $x$ chips of type $g$ in exchange for $y$ chips of type $r$ – they choose a trade to maximize their expected payoff:
Figure 3. Trading patterns in the 3-chip game. Top: Net surplus change for the proposer. Values to the right of the dashed line represent positive surplus to the proposer. Bottom: Chip trading ratio (number of chips offered to requested). Accepted trades are shown in green, trades that were not accepted (for any reason) are in red. Values near the dashed line mark balanced exchange (a 1:1 ratio of chips).
Trading patterns in the 3-chip game. Top: Net surplus change for the proposer. Values to the right of the dashed line represent positive surplus to the proposer. Bottom: Chip trading ratio (number of chips offered to requested). Accepted trades are shown in green, trades that were not accepted (for any reason) are in red. Values near the dashed line mark balanced exchange (a 1:1 ratio of chips).
@@
$\max_{(\mathbf{x},\mathbf{y})}\sum_{\mathbf{v}_{-i}}\mathbf{1}\Bigl\{\exists\,j\neq i:\text{accept}\bigl(v_{j},x_{g},y_{r}\bigr)\Bigr\}\;\times\;\Delta u_{i}(v_{i},x_{g},y_{r})\;\times\;B_{i}\bigl(\mathbf{v}_{-i}\bigr),$
@@
where $\text{accept}\bigl(v_{j},x_{g},y_{r}\bigr)$ indicates that agent $j$ with valuation $v_{j}$ would accept the trade, $\Delta u_{i}(v_{i},x_{g},y_{r})=v_{i,r}y_{r}-v_{i,g}x_{g}$ is agent $i$’s change in utility and $B_{i}(\mathbf{v}_{-i})$ is player $i$’s belief (probability) that the other agents’ valuations are $\mathbf{v}_{-i}$. This agent assumes that their opponent’s acceptance decision is myopically rational, i.e., a receiver with valuation $v_{j}$ accepts if and only if they have sufficient chips for the trade and their expected utility is positive, i.e., $v_{j,g}x_{g}-v_{j,r}y_{r}>0$.
@@
All agents update their priors after observing acceptance decisions; they discard valuation states $\{v_{j}\}$ that contradict the observed event, assuming myopically rational opponents, and re-normalize the remaining probabilities. When the trade is accepted, the proposer keeps only valuations that would be accepted in its distribution over the chosen receiver. The receiver keeps only valuations that would be accepted in its distribution over the proposer. Bystanders update both priors accordingly. When there is no trade, all players discard valuations that conflict with the observed rejections.
6. Results
We organize our findings around the two research questions. For RQ1 (performance), we compare aggregate surplus across agent types. For RQ2 (process), we examine trading patterns both at the turn- and game-level. We additionally report qualitative feedback from human participants on their strategies and experiences.
6.1. Performance
Figure 2 shows how total group surplus—the sum of all players’ gains from trade—evolves over each game’s nine turns. Bayesian agents consistently achieved the highest surplus across all game variants, reaching 74–80% of the optimum. GPT-4o performed comparably to human participants, while Gemini 1.5 Pro achieved lower aggregate surplus (42–45% vs. 54–69% for humans and GPT-4o). The difference in performance between GPT-4o and the humans is negligible. The refined prompting variants did not substantially improve performance for either LLM family.
Beyond the final surplus, several patterns stand out in these trajectories. Bayesian agents show notably concave paths—rapid early gains that flatten as obvious trades are exhausted. In contrast, human and LLM trajectories rise more linearly, with Gemini 1.5 Pro showing the slowest average gains. As game complexity increased from two to four chip types, all agent types captured a smaller share of the optimal surplus, though Bayesian agents remained the most robust to this increase.
6.2. Procedural Alignment
Given that humans and GPT-4o agents achieved comparable surplus, we now ask whether they arrived there through similar processes. In this section, we evaluate strategic differences between the human, GPT-4o, and Bayesian agents. We analyze behavior at two levels, since agents may accept short-term losses to position themselves for later gains, or conversely, capture early value at the cost of foreclosing better opportunities. We first examine the structure of individual proposals at the turn level—what each trade would yield if accepted, without considering future opportunities. We then evaluate decisions retrospectively—whether they were optimal given the opportunities that arose later in the game.
6.2.1. Turn-level analysis
To characterize trading behavior at the turn level, we examine two features of each proposal: (i) the net surplus it would generate for the proposer, and (ii) the ratio of chips offered to chips requested.
Figure 3 shows these distributions for the 3-chip game.1010102- and 4-chip results are qualitatively similar; figures provided in Appendix F.2. Figure values are provided in Table 5. The top row shows that humans generally proposed trades that would generate positive surplus, though some yield net-zero or negative returns. The bottom row shows that human proposals tend to be fair in the number of chips exchanged. Both LLM families tend to propose trades that generate little surplus and more often result in a personal loss. On the chip trading ratio plot, a noticeable tail extends above the 1:1 line, indicating a tendency to offer more chips than requested (ratios up to 5:1).
By construction, Bayesian agents propose trades that give them surplus greater than zero—all density lies right of zero in the surplus plot. [^11] Density mass below zero surplus for the Bayesian agents is an artifact of Gaussian smoothing. None of the Bayesian proposals generate a negative surplus by construction. In the chip trading ratio row, Bayesian proposals cluster below 1:1 (ratio < 1); that is, they request more chips than they offer. Such proposals are frequently rejected.
@@ 6.2.1. Turn-level analysis
These trading patterns reveal clear strategic differences. Humans offer trades suggesting a consideration for balanced, fair exchanges (Danz et al., 2022). LLMs propose highly acceptable trades consistent with a “concessionary posture”—offering many chips in return for a few. Bayesian agents adopt a value-extractive strategy (Fisher et al., 2011), asking to receive more than they give, at the cost of frequent rejections. To evaluate whether any of these strategies were optimal ex-post, we next apply a regret-minimization framework.